OpenStudy (anonymous):

factor -6x^3-3x^2+45x

OpenStudy (anonymous):

Factor out the greatest common factor of all those terms first... what is it?

OpenStudy (anonymous):

x=0,x+3,2x-5

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

-3x(2x^2 + x - 15) = 0 -3x (2x - 5)(x + 3) = 0 x = 0 OR x = 5/2 OR x = -3